# Module schemes in invariant theory

### v1

@@ 17. Fiberflat bundles
Every vector bundle is fiberflat. The common dimension of the fibers is called the rank of the fiberflat bundle. The term fiberflat points to the fact that for a flat family, the dimension of the nonempty fibers is constant, see [26, Proposition 9.5].
@@ 18. Reflexive bundles and factorial closure
Let the finite group $G$ act linearly and faithfully on affine space ${\mathbb{A}}^{d}$ over $K$ with quotient scheme $X=\operatorname{Spec}R$. Let $U\subseteq X$ be the regular locus. Then the following hold.
@@
    The tangent bundle $T|_{U}$ is fiberflat.
@@
The corresponding modules on $S=K[X_{1},\ldots,X_{d}]$ are free of finite rank, with linear representations induced by the basic representation. Their quotient schemes are fiberflat by Lemma 17.2 and they give fiberflat realizations of the objects in question. ∎
@@
The algebra $\bigoplus_{n\in{\mathbb{N}}}(\operatorname{Sym}^{n}M)^{**}$ is called the factorial closure of $M$, see [46, Chapter 7], [38]. The factorial closure gives the canonical fiberflat realization of a reflexive fiberflat module.
@@
is finite by [24, Corollaire 5.11.4], it is an isomorphism above $U$, thus birational, and therefore, it is the normalization, as $V$ is normal. We also have
@@ 19. Fiberflat bundles on regular rings
Let $R$ be a noetherian excellent local domain of dimension $d\geq 3$, $M$ a finitely generated $R$-module of rank $r$. Let $U\subseteq\operatorname{Spec}R$ denote the locus where $M$ is locally free, suppose that $U$ is regular, contains all points of codimension one, and set $Y=X\setminus U$. Let $p:Z\rightarrow\operatorname{Spec}R$ be an irreducible geometric realization of $M$, which is not fiberflat. Suppose that the dimension of $p^{-1}(Y)$ is $\leq r+d-2$. Then, $M$ is not fiberflat.
@@
The dimension of $Z$ is $\geq d+r$ (look at a chain of prime ideals on the zero section and then continue generically), so the complement of the vector bundle $p^{-1}(U)$ in $Z$ has codimension $\geq 2$. Hence, $A\rightarrow\Gamma(p^{-1}(U),{\mathcal{O}}_{Z})=C$ is a finite homomorphism by [24, Corollaire 5.11.4] and $C$ is the normalization of $A$. Then $C$ gives a realization which is also not fiberflat. Suppose that $q:\operatorname{Spec}B\rightarrow X$ is an integral fiberflat geometric realization of $M$. Then the dimension of $q^{-1}(Y)$ is $\leq r+d-2$, since $\dim Y\leq d-2$. Therefore, $C$ is also the normalization of $B$, which yields a contradiction. ∎
@@
Let $(R,{\mathfrak{m}})$ be noetherian excellent local domain of dimension $d\geq 3$, $M$ a finitely generated $R$-module such that $M$ is locally free outside ${\mathfrak{m}}$. Suppose that $M$ is fiberflat and $\nu(M)\leq r+d-2$. Then, $M$ is free.
@@
Let $R=K[X_{1},\ldots,X_{d}]^{G}$ be an invariant ring for a linear action of a finite group. Suppose that $R$ has an isolated singularity and that $\operatorname{Spec}(\operatorname{Sym}(\Omega_{R|K}))$ is irreducible. Then the embedding dimension of $R$ is $\geq 2d-1$.
@@
We suspect that for a regular ring, Corollary 19.3 holds without the condition on the number of generators. The intuition behind this is the following: A fiberflat bundle has the property that going modulo a regular element of the base ring, the constance of the dimension is preserved. By Lemma 17.5, a fiberflat bundle is irreducible. Hence, a fiberflat bundle behaves in some geometric sense like a maximal Cohen-Macaulay module. And for a regular ring, a maximal Cohen-Macaulay module is free.
@@
We consider $S=K[X,Y,Z]$ with the action of ${\mathbb{Z}}/(2)$ by negation on every variable. The invariant ring $R$ is the second Veronese algebra of dimension three, which is the only invariant ring of dimension $\geq 3$ with only finitely many maximal Cohen-Macaulay modules, see [4, Theorem 4.1], [30, Theorem 16.5]. We consider, according to Lemma 19.9, the exact sequence
@@ 21. Singularities
We cite the following fact, which is a generalization of the theorem of Chevalley, Shephard, Todd and Serre, as a reference, we only found [19, Satz 1] in the complex case (see also [36]). Therefore, we will restrict to $K={\mathbb{C}}$ in this section, but we will continue to write $K$.
@@
Let $K={\mathbb{C}}$. Let $\beta$ be a linear faithful representation of a finite group $G$, let $X={\mathbb{A}}^{d}/G$ be the quotient and let $Q\in{\mathbb{A}}^{d}$ be a closed point with image point $P\in X$. Let $\operatorname{Stab}(Q)$ be the stabilizer of $Q$. Then, $P$ is a regular point if and only if $\beta(\operatorname{Stab}(Q))\subseteq\operatorname{Gl}(K^{d})$ is generated by reflections.
The following fact is known as theorem of Steinberg, see [43, Theorem 1.5], we include a short proof based on the previous lemma.
@@
Let $G\subseteq\operatorname{Gl}_{d}({\mathbb{C}})$ be a reflection group and let $Q\in{\mathbb{A}}^{d}$ be a point. Then the stabilizer group $\operatorname{Stab}(Q)$ is also a reflection group.
@@ 22. The fundamental group of quotient schemes
The quotient $={\mathbb{A}}^{d}/G_{\beta\operatorname{refl}}$ is isomorphic to the affine space, $\overline{G}$ is acting on it and the quotient of this action is also $X$, as mentioned in the beginning of Section 15. So we can assume that the action $\beta$ is small. Let $V\subseteq{\mathbb{A}}^{d}$ be the free locus and let $U=V/G$ be its image. The complement ${\mathbb{A}}^{d}\setminus V$ has codimension $\geq 2$ since the group $G$ contains no reflection, and hence, $X\setminus U$ also has codimension $\geq 2$. The mapping $V\rightarrow U$ is a covering space, $U$ is a complex manifold, $V$ is simply connected, as ${\mathbb{A}}^{d}$ is and since the fundamental group does not change by removing closed real submanifolds of real codimension $\geq 3$, by [44, Satz 4.B.2] (applied to a smooth stratification of ${\mathbb{A}}^{d}\setminus V$). Therefore, $G$ is isomorphic to the fundamental group of $U$. By Corollary 21.4, $U$ is the regular locus of $X$. ∎
@@
    We have $\pi_{1}(Z_{\rho\operatorname{reg}}\setminus p^{-1}(X_{\operatorname{reg}}))=\pi_{1}(Z_{\rho\operatorname{reg}})$.
@@
    In the case of a small basic action $\beta$, the open subset $Z_{\rho\operatorname{reg}}\setminus p^{-1}(X_{\operatorname{reg}})$ coincides with the vector bundle $Z_{\rho}|_{U}$, where $U\subseteq X$ is the image of the fixed point free locus, and its fundamental group is $G$.
@@
(1). The singular locus of $X$ has codimension $\geq 2$, as $X$ is normal, and since the fibers of $Z_{\rho}$ have the same dimension as noted in Remark 6.3, $p^{-1}(X_{\operatorname{reg}})$ also has codimension at least $2$. Hence, $\pi_{1}(Z_{\rho\operatorname{reg}})\cong\pi_{1}(Z_{\rho\operatorname{reg}}\setminus p^{-1}(X_{\operatorname{reg}}))$ by the argument applied in the proof of Lemma 22.1.
(2). In the case of a small basic, $U$ is the regular locus of $X$ by Corollary 21.4. $Z_{\rho}|_{U}$ is a vector bundle by Corollary 5.14, hence $Z_{\rho})_{U}=Z_{\rho\operatorname{reg}}\setminus p^{-1}(X_{\operatorname{reg}})$. The claim about the fundamental group follows from Corollary 22.8. ∎
@@
$\pi_{1}(Z_{\rho\operatorname{reg}})\cong\pi_{1}(Z_{\rho\operatorname{reg}}\setminus p^{-1}(X_{\operatorname{reg}}))\longrightarrow\pi_{1}(X_{\operatorname{reg}})$
@@
On the left, $V$ is the fixed point free locus of the action of the small group $G/H$ on ${\mathbb{A}}^{d}$, $V\rightarrow U$ is a covering space, $V$ is simply connected and $U=X_{\operatorname{reg}}$. On the right, $\tilde{V}$ is the fixed point free locus of the product action of $G$ on ${\mathbb{A}}^{d}\times{\mathbb{A}}^{m}$, whose complement has codimension $\geq 2$ since $\beta\times\rho$ is small. Hence, $\tilde{V}$ has trivial fundamental group. On the right, we have covering spaces, and $\tilde{V}/G=Z_{\rho\operatorname{reg}}$. On the very right, $\hat{V}$ is such that $\hat{V}/G=Z_{\rho\operatorname{reg}}\setminus p^{-1}(X_{\operatorname{reg}})$, which does not change the fundamental groups. The horizontal morphisms map $\hat{V}/G$ to $U$ and $\hat{V}/H$ to $V$.
We fix a point $R\in Z_{\rho\operatorname{reg}}\setminus p^{-1}(X_{\operatorname{reg}})=\hat{V}/G$ with image point $P\in U$ and we fix a point $T\in\hat{V}$ above $R$. An element $\gamma\in\pi_{1}(Z_{\rho\operatorname{reg}}\setminus p^{-1}(X_{\operatorname{reg}}))$ with starting point $R$ corresponds (see Remark 22.2) to a lifting $\hat{\gamma}$ in $\hat{V}$ starting in $T$ and ending in $T^{\prime}$, and $\gamma$ corresponds to the unique element $g\in G$ with $g(T)=T^{\prime}$. Let $S$ and $S^{\prime}$ be the image points of $T$ and $T^{\prime}$ in $\hat{V}/H$. Then, we have the relation $[g](S)=S^{\prime}$ for $[g]\in G/H$. Let $Q$ and $Q^{\prime}$ be the image points of $S$ and $S^{\prime}$ in $V$. There, we also have the relation $[g](Q)=Q^{\prime}$. The corresponding relations hold for the various images of the path $\hat{\gamma}$ in $\hat{V}/H$, in $V$ and in $U$, i.e., $S^{\prime}$ and $Q^{\prime}$ are the endpoints of the liftings. Therefore, the image of $\gamma$ in $\pi_{1}(U)$ corresponds to $[g]$. ∎
@@
We continue with Example 21.7. The fundamental group of the punctured spectrum of the quotient scheme $X={\mathbb{A}}^{2}/G\cong{\mathbb{A}}^{2}/(G/H)$ is $G/H\cong{\mathbb{Z}}/(2)$ according to Lemma 22.1. The singular locus of $Z_{\rho}$ is the line $V(X_{2},W)$ mapping to a curve in $X$, and its complement $Z_{\rho\operatorname{reg}}$ has fundamental group $G$ according to Corollary 22.7. The set $Z_{\rho\operatorname{reg}}\setminus p^{-1}(X_{\operatorname{reg}})$ is the complement of two curves meeting in a point, its fundamental group is also $G$, and the natural mapping between the fundamental groups coincides with $G\rightarrow G/H$ by Theorem 22.10.

### v2

@@ 17. Fiberflat bundles
Every vector bundle is fiberflat. The common dimension of the fibers is called the rank of the fiberflat bundle. The term fiberflat points to the fact that for a flat family, the dimension of the nonempty fibers is constant, see [28, Proposition 9.5].
@@
Lemma 17.7.
Let $R$ be a domain of finite type over a field $K$ and let $p:Z\rightarrow\operatorname{Spec}R$ be an affine scheme of finite type which is locally free over $U=\operatorname{Spec}R\setminus\{{\mathfrak{m}}\}$, where ${\mathfrak{m}}$ denotes a maximal ideal of ${\mathfrak{m}}$. Suppose that $p^{-1}({\mathfrak{m}})$ is irreducible. Then, $Z$ is fiberflat if and only if the morphism is open.
Proof.
We use the going down criteria for openness, see [25, Corollaire 1.10.4]. If $Z\rightarrow\operatorname{Spec}R$ is not fiberflat, then for an irreducible curve $C$ through ${\mathfrak{m}}$ the restriction $Z|_{C}$ is not irreducible. A point above ${\mathfrak{m}}$ which lives in the component above ${\mathfrak{m}}$ but not in the dominant component does not live on a curve dominating $C$.
Suppose now that $Z\rightarrow\operatorname{Spec}R$ is fiberflat, and let ${\mathfrak{p}}\subseteq{\mathfrak{q}}$ be a chain of prime ideals in $R$ and ${\mathfrak{q}}^{\prime}$ a prime ideal in $Z$ mapping to ${\mathfrak{q}}$. We look at the situation above $V({\mathfrak{p}})$, the fiberflattness is preserved. By Lemma 17.5, $Z|_{V({\mathfrak{p}})}$ is irreducible, ${\mathfrak{q}}^{\prime}$ lies on it and its generic point is mapped to ${\mathfrak{p}}$. ∎
@@ 18. Reflexive bundles and factorial closure
Let the finite nonmodular group $G$ act linearly and faithfully on affine space ${\mathbb{A}}^{d}$ over $K$ with quotient scheme $X=\operatorname{Spec}R$. Let $U\subseteq X$ be the regular locus. Then the following hold.
@@
    The tangent bundle $T_{U}$ is fiberflat.
@@
We mod out the action of the subgroup generated by the reflections and can then assume that the action $\beta$ is small. The action is then free on a subset $V$ containing all points of codimension one with image $U\subseteq X$, which is the regular locus of $X$. For the tangent bundle on $U$, we have by Corollary 5.15, $T_{U}=T_{V}/\beta\times\beta=V\times{\mathbb{A}}^{d}/\beta\times\beta$. The quotient scheme $T_{{\mathbb{A}}^{d}}/\beta\times\beta={\mathbb{A}}^{d}\times{\mathbb{A}}^{d}/\beta\times\beta$ gives by Lemma 17.2 a fiberflat realization of $T_{U}$. The proof of th second statement is analogous. ∎
@@
The algebra $\bigoplus_{n\in{\mathbb{N}}}(\operatorname{Sym}^{n}M)^{**}$ is called the factorial closure of $M$, see [50, Chapter 7], [42]. The factorial closure gives the canonical fiberflat realization of a reflexive fiberflat module.
@@
is finite by [26, Corollaire 5.11.4], it is an isomorphism above $U$, thus birational, and therefore, it is the normalization, as $V$ is normal. We also have
@@ 19. Fiberflat bundles on regular rings
Let $R$ be a noetherian excellent local domain of dimension $d\geq 3$, $M$ a finitely generated $R$-module of rank $r$. Let $U\subseteq\operatorname{Spec}R$ denote the locus where $M$ is locally free, suppose that $U$ is regular, contains all points of codimension one, and set $Y=X\setminus U$. Let $p:Z\rightarrow\operatorname{Spec}R$ be an irreducible surjective geometric realization of $M$, which is not fiberflat. Suppose that the dimension of $p^{-1}(Y)$ is $\leq r+d-2$. Then, $M$ is not fiberflat.
@@
The dimension of $Z$ is $\geq d+r$ (look at a chain of prime ideals on the zero section and then continue generically), so the complement of the vector bundle $p^{-1}(U)$ in $Z$ has codimension $\geq 2$. Hence, $A\rightarrow\Gamma(p^{-1}(U),{\mathcal{O}}_{Z})=C$ is a finite homomorphism by [26, Corollaire 5.11.4] and $C$ is the normalization of $A$. Then $C$ gives a normal realization which is also not fiberflat. Suppose that $q:\operatorname{Spec}B\rightarrow X$ is an integral fiberflat geometric realization of $M$. Then the dimension of $q^{-1}(Y)$ is $\leq r+d-2$, since $\dim Y\leq d-2$. Therefore, $C$ is also the normalization of $B$, which yields a contradiction. ∎
@@
Example 19.3.
Let $R=K[X,Y,Z]$ and consider the module $M=R^{3}/(Xe_{2},Ye_{2},Ze_{2},Xe_{3},Ye_{3},Ze_{3})$, its symmetric algebra is
$R[A,B,C]/(XB,YB,ZB,XC,YC,ZC)\,$
which is reducible. $M$ is fiberflat, as on the punctured spectrum, it is just the free module of rank one, the dimension of the fiber above $(X,Y,Z)$ is $3$. This shows that we need the irreducibility condition in Corollary 19.2. In the situation of this corollary, one can try to apply Lemma 19.1 to the symmetric algebra modulo torsion, but it may happen, as in the example, that the exceptional dimension does not contradict fiberflatness anymore.
@@
Let $(R,{\mathfrak{m}})$ be noetherian excellent local domain of dimension $d\geq 3$, $M$ a finitely generated $R$-module such that $M$ is locally free outside ${\mathfrak{m}}$. Suppose that $M$ is fiberflat, that the symmetric algebra $\operatorname{Sym}M$ is irreducible and that $\nu(M)\leq r+d-2$. Then, $M$ is free.
@@
Let $R=K[X_{1},\ldots,X_{d}]^{G}$ be an invariant ring for a linear action of a finite nonmodular group. Suppose that $R$ has an isolated singularity and that the symmetric algebra $\operatorname{Spec}(\operatorname{Sym}(\Omega_{R|K}))$ is irreducible. Then the embedding dimension of $R$ is $\geq 2d-1$.
@@
We suspect that for a regular ring, Corollary 19.4 holds without the condition on the number of generators. The intuition behind this is the following: A fiberflat bundle has the property that going modulo a regular element of the base ring, the constance of the dimension is preserved. By Lemma 17.5, a fiberflat bundle with an irreducible central fiber is irreducible. Hence, a fiberflat bundle behaves in some geometric sense like a maximal Cohen-Macaulay module. And for a regular ring, a maximal Cohen-Macaulay module is free.
Remark 19.7.
By [19, Theorem 1.1], the flatness of $R\rightarrow A$ for $R$ regular, everything of finite type over ${\mathbb{C}}$, is equivalent to the property that $A\otimes_{R}A\otimes_{R}\ldots\otimes_{R}A$ ($\dim R$ factors) is torsionfree (hint of Adrian Langer). Fiberflatness implies at least that this tensor product is irreducible, as the argument in Lemma 17.5 shows.
@@
For the following example, I thank Craig Huneke for a hint.
@@
We consider $S=K[X,Y,Z]$ with the action of ${\mathbb{Z}}/(2)$ by negation on every variable. The invariant ring $R$ is the second Veronese algebra of dimension three, which is the only invariant ring of dimension $\geq 3$ with only finitely many maximal Cohen-Macaulay modules, see [4, Theorem 4.1], [32, Theorem 16.5]. We consider, according to Lemma 19.11, the exact sequence
@@ 21. Singularities
We use the following fact, which is a slight generalization of the theorem of Chevalley, Shephard, Todd and Serre, as a reference, we only found [20, Satz 1] in the complex case (see also [39]). I thank Mitsuyasu Hashimoto for help with the following proof.
@@
Let $\beta$ be a linear faithful representation of a nonmodular finite group $G$, let $X={\mathbb{A}}^{d}/G$ be the quotient and let $Q\in{\mathbb{A}}^{d}$ be a closed point with image point $P\in X$. Let $\operatorname{Stab}(Q)$ be the stabilizer of $Q$. Then, $P$ is a regular point if and only if $\beta(\operatorname{Stab}(Q))\subseteq\operatorname{Gl}(K^{d})$ is generated by reflections.
Proof.
We may assume that $K$ is algebraically closed. Let $H=\operatorname{Stab}(Q)$, $Q=(a_{1},\dots,a_{d})$, $Y_{i}=X_{i}-a_{i}$. Let ${\mathfrak{q}}\subseteq S=K[X_{1},\ldots,X_{d}]$ be the maximal ideal corresponding to $Q$, ${\mathfrak{p}}={\mathfrak{q}}\cap S^{G}$. From the ring homomorphisms $S^{G}\rightarrow S^{H}\rightarrow S$ we get natural ring homomorphisms
$(S^{G})_{\mathfrak{p}}\longrightarrow(S^{H})_{{\mathfrak{q}}\cap S^{H}}\cong(S_{\mathfrak{q}})^{H}\longrightarrow S_{\mathfrak{q}}\,.$
The homomorphism $(S^{G})_{\mathfrak{p}}\rightarrow(S_{\mathfrak{q}})^{H}$ is unramified by [37, Theorem 4.1.2], hence, $(S^{G})_{\mathfrak{p}}$ is regular if and only if $(S_{\mathfrak{q}})^{H}$ is regular. But we can understand the action of $H$ on ${\mathbb{A}}^{d}$ as a linear action with respect to the new variables $Y_{i}$, and then apply the Theorem of Chevalley, Shephard, Todd. ∎
The following fact is known as theorem of Steinberg, see [47, Theorem 1.5], we include a short proof based on the previous lemma.
@@
Let $G\subseteq\operatorname{Gl}_{d}(K)$ be a nonmodular reflection group and let $Q\in{\mathbb{A}}^{d}$ be a point. Then the stabilizer group $\operatorname{Stab}(Q)$ is also a reflection group.
@@ 22. The fundamental group of quotient schemes
The quotient $={\mathbb{A}}^{d}/G_{\beta\operatorname{refl}}$ is isomorphic to the affine space, $\overline{G}$ is acting on it and the quotient of this action is also $X$, as mentioned in the beginning of Section 15. So we can assume that the action $\beta$ is small. Let $V\subseteq{\mathbb{A}}^{d}$ be the free locus and let $U=V/G$ be its image. The complement ${\mathbb{A}}^{d}\setminus V$ has codimension $\geq 2$ since the group $G$ contains no reflection, and hence, $X\setminus U$ also has codimension $\geq 2$. The mapping $V\rightarrow U$ is a covering space, $U$ is a complex manifold, $V$ is simply connected, as ${\mathbb{A}}^{d}$ is and since the fundamental group does not change by removing closed real submanifolds of real codimension $\geq 3$, by [48, Satz 4.B.2] (applied to a smooth stratification of ${\mathbb{A}}^{d}\setminus V$). Therefore, $G$ is isomorphic to the fundamental group of $U$. By Corollary 21.4, $U$ is the regular locus of $X$. ∎
@@
    We have $\pi_{1}(Z_{\rho\operatorname{reg}}\cap p^{-1}(X_{\operatorname{reg}}))=\pi_{1}(Z_{\rho\operatorname{reg}})$.
@@
    In the case of a small basic action $\beta$, the open subset $Z_{\rho\operatorname{reg}}\cap p^{-1}(X_{\operatorname{reg}})$ coincides with the vector bundle $Z_{\rho}|_{U}$, where $U\subseteq X$ is the image of the fixed point free locus, and its fundamental group is $G$.
@@
(1). The singular locus of $X$ has codimension $\geq 2$, as $X$ is normal, and since the fibers of $Z_{\rho}$ have the same dimension as noted in Remark 6.3, $p^{-1}(X_{\operatorname{reg}})$ also has codimension at least $2$. Hence, $\pi_{1}(Z_{\rho\operatorname{reg}})\cong\pi_{1}(Z_{\rho\operatorname{reg}}\cap p^{-1}(X_{\operatorname{reg}}))$ by the argument applied in the proof of Lemma 22.1.
(2). In the case of a small basic, $U$ is the regular locus of $X$ by Corollary 21.4. $Z_{\rho}|_{U}$ is a vector bundle by Corollary 5.14, hence $Z_{\rho})_{U}=Z_{\rho\operatorname{reg}}\cap p^{-1}(X_{\operatorname{reg}})$. The claim about the fundamental group follows from Corollary 22.8. ∎
@@
$\pi_{1}(Z_{\rho\operatorname{reg}})\cong\pi_{1}(Z_{\rho\operatorname{reg}}\cap p^{-1}(X_{\operatorname{reg}}))\longrightarrow\pi_{1}(X_{\operatorname{reg}})$
@@
On the left, $V$ is the fixed point free locus of the action of the small group $G/H$ on ${\mathbb{A}}^{d}$, $V\rightarrow U$ is a covering space, $V$ is simply connected and $U=X_{\operatorname{reg}}$. On the right, $\tilde{V}$ is the fixed point free locus of the product action of $G$ on ${\mathbb{A}}^{d}\times{\mathbb{A}}^{m}$, whose complement has codimension $\geq 2$ since $\beta\times\rho$ is small. Hence, $\tilde{V}$ has trivial fundamental group. On the right, we have covering spaces, and $\tilde{V}/G=Z_{\rho\operatorname{reg}}$. On the very right, $\hat{V}$ is such that $\hat{V}/G=Z_{\rho\operatorname{reg}}\cap p^{-1}(X_{\operatorname{reg}})$, which does not change the fundamental groups. The horizontal morphisms map $\hat{V}/G$ to $U$ and $\hat{V}/H$ to $V$.
We fix a point $R\in Z_{\rho\operatorname{reg}}\cap p^{-1}(X_{\operatorname{reg}})=\hat{V}/G$ with image point $P\in U$ and we fix a point $T\in\hat{V}$ above $R$. An element $\gamma\in\pi_{1}(Z_{\rho\operatorname{reg}}\cap p^{-1}(X_{\operatorname{reg}}))$ with starting point $R$ corresponds (see Remark 22.2) to a lifting $\hat{\gamma}$ in $\hat{V}$ starting in $T$ and ending in $T^{\prime}$, and $\gamma$ corresponds to the unique element $g\in G$ with $g(T)=T^{\prime}$. Let $S$ and $S^{\prime}$ be the image points of $T$ and $T^{\prime}$ in $\hat{V}/H$. Then, we have the relation $[g](S)=S^{\prime}$ for $[g]\in G/H$. Let $Q$ and $Q^{\prime}$ be the image points of $S$ and $S^{\prime}$ in $V$. There, we also have the relation $[g](Q)=Q^{\prime}$. The corresponding relations hold for the various images of the path $\hat{\gamma}$ in $\hat{V}/H$, in $V$ and in $U$, i.e., $S^{\prime}$ and $Q^{\prime}$ are the endpoints of the liftings. Therefore, the image of $\gamma$ in $\pi_{1}(U)$ corresponds to $[g]$. ∎
@@
We continue with Example 21.7. The fundamental group of the punctured spectrum of the quotient scheme $X={\mathbb{A}}^{2}/G\cong{\mathbb{A}}^{2}/(G/H)$ is $G/H\cong{\mathbb{Z}}/(2)$ according to Lemma 22.1. The singular locus of $Z_{\rho}$ is the line $V(X_{2},W)$ mapping to a curve in $X$, and its complement $Z_{\rho\operatorname{reg}}$ has fundamental group $G$ according to Corollary 22.7. The set $Z_{\rho\operatorname{reg}}\cap p^{-1}(X_{\operatorname{reg}})$ is the complement of two curves meeting in a point, its fundamental group is also $G$, and the natural mapping between the fundamental groups coincides with $G\rightarrow G/H$ by Theorem 22.10.